OpenStudy (anonymous):

What is radical 216k^4 show work

OpenStudy (anonymous):

Just divided power on 2 :)

OpenStudy (anonymous):

Wait what?

OpenStudy (anonymous):

Look at this example : \[\sqrt{x^6}=x^6/2\] Got it ?:)

OpenStudy (anonymous):

Yes thanks canu pleases show your work??(:

OpenStudy (anonymous):

In this question ?

OpenStudy (anonymous):

For the question I asked please

OpenStudy (anonymous):

K : \[\sqrt{216k^4}=216k^4/2=216k^2\] Got it ?:)

OpenStudy (anonymous):

Okay so the answer is 216k^2??

OpenStudy (anonymous):

Yep !

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

You re welcome !

OpenStudy (anonymous):

Does their need to be any radicals in the answer?

OpenStudy (anonymous):

What ???

OpenStudy (anonymous):

Any symbol! Like square root

OpenStudy (anonymous):

K ... You want example ?

OpenStudy (anonymous):

No I just want to know if in the answer do I need to put the square root symbol in the answer

OpenStudy (anonymous):

Yes !

OpenStudy (anonymous):

@Sophieeee